CARS AND INFRARED STUDIES OF THE ν_{1}, ν_{2} AND ν_{4} BANDS OF ^{34}S^{18}O_{3}

\underline{TONY MASIELLO},  JEFFREY  BARBER,  ENGELENE  t.H.  CHRYSOSTOM,

JOSEPH W. NIBLER , Department of Chemistry, Oregon State University, Corvallis, OR 97331; ARTHUR MAKI, 15012 24th Ave. S. E. Mill Creek, WA 98012; ALFONS WEBER, National Science Foundation, Arlington, VA 22230 and National Institute of Standards and Technology, Gaithersburg, MD 20899; THOMAS A. BLAKE, ROBERT L. SAMS, Pacific Northwest National Laboratory, Richland, WA 99352.

We are engaged in a comprehensive investigation of the spectroscopic properties of sulfur trioxide, an important participant in reactions in the upper atmosphere. The fundamental modes and several hot bands of the isotopic variants (^{32}S^{18}O_{3}, 34S^{16}O_{3}, and ^{34}S^{18}O_{3}) have been investigated using high resolution infrared spectroscopy and coherent anti-Stokes Raman scattering. For all isotopic variants, the Raman-active symmetric stretching mode ν_{1} shows complex Q-branch patterns due to indirect Coriolis couplings, l-resonances, and Fermi resonances with dark ν_{2}, ν_{4} combination/overtone levels. Essential to modeling the interactions of these levels with ν_{1} is the understanding of the fundamental vibrations that make up these levels. The analysis of the ν_{2}, ν_{4} infrared active fundamental vibrations of ^{34}S^{18}O_{3} will be presented, along with efforts to model the complex ν_{1} CARS spectrum using information derived from studies of hot bands involving ν_{2} and ν_{4}.